A car accelerates from 0 to 110. km/h in 4.20 s. 1) What force does a 64.0-kg passenger experience during this acceleration? (Express your answer to three significant figures.)
Answer #1

initial velocity = V0 = 0 km/h

final velocity = V = 110 * 5 / 18 = 30.56 m/sec

time = t = 4.2 sec

    V = V0 + a *t

30.56 = 0 + a * 4.2

acceleration = a = 7.27 m/sec^2

force = F = m a = 64 * 7.27 = 465.6 = 466 N